OpenStudy (anonymous):

Solve the Equation 25.14=d + (-15.3)

HanAkoSolo (jamierox4ev3r):

alright are you solving for d in this?

OpenStudy (anonymous):

yeah

HanAkoSolo (jamierox4ev3r):

d-15.3=25.14 +15.3 +15.3 ------------ d=9.84

HanAkoSolo (jamierox4ev3r):

i came up with the final solution that d=9.84. is that a choice? xD

OpenStudy (anonymous):

it must have been -9.84 because that was the other choice apart from 9.84 but no worries thanks for trying. If i need any more help ill be sure to ask you :)

HanAkoSolo (jamierox4ev3r):

are you sure you wrote the equation right? I checked it on the calculator, and it is definitely positive! hmmm :/

OpenStudy (anonymous):

Yeah I mean I also did it on the calculator and got the same answer but for some reason it said it was wrong.

OpenStudy (anonymous):

Wierd huh? Well the good thing is that was the only question I missed :P

OpenStudy (anonymous):

@AdrianGolfWang :May I join and help ?:)

OpenStudy (anonymous):

Sure why not :)

HanAkoSolo (jamierox4ev3r):

this is really weird.. lol *shrugs*

OpenStudy (anonymous):

haha tell me about it, Dont worry about it im over that question :)

OpenStudy (anonymous):

\[Find The Difference 11\frac{ 9 }{ 14 } - 9\frac{ 3 }{ 7 }\]

HanAkoSolo (jamierox4ev3r):

well just make sure that its 11\(\frac{9}{14}\) and -9\(\frac{6}{14}\) so that they are the same denominator.

HanAkoSolo (jamierox4ev3r):

so 11-9 is 2, and 9/14 - 6/14 = 3/14... so your final answer is 2\(\frac{3}{14}\). Make sense @AdrianGolfWang ?

OpenStudy (anonymous):

Thanks Jamie you rox 4 ever! lol

OpenStudy (anonymous):

Yeah...got it Ty :)